If 0.485 g of K_2CO_3 were reacted according to th
Answer #1

Given

mass of K2CO3 = 0.485g

number of moles of K2CO3 = Weight / Molar mass = 0.485g / 138.205 g/mol = 0.0035

from the reaction,

1 mole of K2CO3 gives 2 moles of KCl

then 0.0035 moles of K2CO3 gives (2 * 0.00350) = 0.00701 moles of KCl

therefore weigth of KCl formed = number of moles * molar mass of KCl = 0.00701 * 74.5513 = 0.5232gr.